OpenStudy (anonymous):

a packet of 3 greeting cards at a shop was originally priced at 7.95. During a sale, the price was reduced by 20%. What was the price of each greeting card in the packet during the sale?

OpenStudy (anonymous):

\[7.95\frac{20}{100} = 1.59\] gets 20% of the price. Taking it away: \[7.95 - 1.59 = 6.36\] Dividing the answer by 3 to get the price of each card: \[\frac{6.36}{3}=1.12\] so the answer is 2.12